OpenStudy (anonymous):

Alright, last one! 5/2x+6 = 1/4x ?! I somehow got 13/5 but that is definitely not correct

OpenStudy (phi):

can you be more clear about the equation. Use the equation editor, or put in parens, because it's not clear what is in the denominator and what is in the numerator.

OpenStudy (phi):

\[ \frac{5}{2}x+6 = \frac{1}{4} x \]?

OpenStudy (anonymous):

Sorry! I should have closed it! I figured it out! But just for your quench your curiosity, it was: 5/(2x+6) = 1/(4x)

OpenStudy (anonymous):



OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

The answer I got was, 1/3

OpenStudy (phi):

yes, and the easiest way (doing it by hand i.e. not wolfram) is cross multiply to get 5*4x = 2x+6 20x = 2x + 6

OpenStudy (anonymous):

Thank you!